A radiator contains 9 quarts of fluid, 30% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 40% antifreeze? (Round your answer to two decimal places.)
Answer by richwmiller(17219)

0.3*9-0.3x+1x=0.4(9)
2.7+0.7x=3.6
0.7x=3.6-2.7
0.7x=0.9
x=1.28571429
check
0.3*7.71428571+1*1.28571429=0.4(9)
2.31428571+1.28571429=3.6
3.6=3.6
ok
codemixrep